# BIV-Priv-Seg: Locating Private Content in Images Taken by People With Visual Impairments

### v2

@@ 4.2 VLM Localization
We next benchmarked the top-performing, off-the-shelf vision language model that supports visual grounding in a zero-shot settings: GLaMM [39]. We explored prompting the model in the following three ways:
Prompt I. Privacy-agnostic (category specific): Each category’s name is inserted into the following prompt: “Segment it if there is a {category name} in the given image, otherwise say there is no {category name}.” This scenario is very general, rather than specialized in the concept of privacy.
Prompt II. Privacy-aware: We use the following generic prompt “Segment it if there is a private object in the given image, otherwise say there is no private object.”, where the specific category name is replaced with the generic phrase “private object”. This prompt captures is intended to capture the model’s ability to understand the concept of privacy.
Prompt III. Privacy-explained: To provide the context of privacy, we provide the list of private categories in our dataset and then request the segmentation as follows: “Private objects include {all category names}. Segment it if there is a private object in the given image, otherwise say there is no private object.”
We intentionally designed these model prompts to allow for a negative response regarding the presence of the target, meaning that the model does not return a segmentation result. We achieved this by extending the grounding prompts from previous work [39] from “Segment something” to “Segment it if something, otherwise say there is no something”.
@@
Privacy classification: We first evaluate the model performance on two levels: (1) Semantic level prediction, which is based on the language used in the conversation, where positive responses such as “yes, here is the mask” or “sure, here is the mask” indicate the presence of a target, and negative responses such as “no, there is no…” indicate the absence of a target; and (2) Mask level prediction, which is determined by whether any pixel in the generated mask has a positive prediction. The results are presented in accuracy, recall, and true negative rate. Accuracy shows the overall performance; recall reflects the percentage of correct predictions among the target-containing images; true negative rate reflects the percentage of correct predictions among the target-less images.
Segmentation: For the images that are classified as private on both semantic and mask levels, we then perform localization. The mIoU and cIoU metrics are used to evaluate the semantic segmentation results across all categories in each scenario. As most of the images in this dataset only contain one target object, we convert those segmentation results into instance segmentation format and evaluate the models in AP50 and Recall metrics.
This two-step evaluation highlights the model’s ability to identify whether there are private objects in the image and its localization ability.
Overall Performance.
Below, we present the results of classification and segmentation, respectively.
The results of both the semantic level and mask level privacy classifications are shown in Table 4. Two key findings emerge. First, on the semantic level, the model consistently generates positive responses across all scenarios, producing masks for target objects in every given image. However, some of these generated masks contain negative results, indicating that while the model confirms the presence of a private object at the conversational level, the masks do not actually predict any pixels corresponding to the private object. This inconsistency suggests that the model lacks a robust text-mask alignment mechanism. Second, despite some negative predictions at the mask level for prompt II and prompt III, the true negative (TN) rates are still extremely low. This indicates the model’s inability to accurately identify when there are no target objects in the image. Prompt I results, which are aggregated across all object categories, show that at least one category is predicted to contain the target in every image in BIV-Priv-Seg. We suspect that both issues stem from biases in the training data, where there is a lack of target-less samples.
@@
Prompt II 69.26 83.97 5.21 81.32 100.00 0.00
Prompt III 63.71 74.28 17.71 81.32 100.00 0.00
Table 4: Classification performance of GLaMM on BIV-Priv-Seg in classifying the images into target-containing and target-less groups on both the level of semantics and masks. Results are presented in accuracy (Acc), recall (R), and true negative rate (TN) as percentages.
Results for the target-containing group are shown in Table 5, with qualitative results in Figure 5. We first compared the performance of GLaMM on our dataset to those on VizWiz-FewShot [47], GranD [39], refCOCO, refCOCO+, and refCOCOg [24]. The results show that the performance of our dataset is comparable to existing datasets. Specifically, the performance on refCOCO, refCOCO+, and refCOCOg is better than BIV-Priv-Seg (prompt I), while the performance on GranD and VizWiz-FewShot is worse. This indicates that state-of-the-art VLMs have the potential to effectively ground private objects in images, achieving performance comparable to other established datasets. We then compared the performance of prompts I, II, and III to explore the model’s ability to comprehend privacy. We observed that prompt I achieved the highest performance, indicating that when given a specific object category, the model has the potential to locate private objects accurately. Conversely, prompt II had the worst performance. Even though the categories in BIV-Priv-Seg are perceived as private by humans and the semantic meaning of prompt II encompasses these categories, the model struggled to localize the private objects as effectively as it did with prompt I. This suggests that the model’s understanding of privacy is less nuanced than human perception. Interestingly, while prompt III attempted to mitigate this semantic ambiguity by providing an explanation of privacy, it still performed poorly compared to prompt I. Despite the initial definition of privacy in prompt III, the model failed to fully comprehend and execute the instructions. This indicates that the model is highly sensitive to the sentence structure of prompts, which may reflect biases in its training data.
Dataset mIoU cIoU AP50 Recall
BIV-Priv-Seg (ours, prompt I) 68.0 51.0 44.7 55.9
BIV-Priv-Seg (ours, prompt II) 11.3 -26.1 3.1 7.8
BIV-Priv-Seg (ours, prompt III) 20.6 -0.5 9.4 14.8
VizWiz-FS [47] (prompt I) 15.6 -16.1 2.2 5.6
@@
Table 5: Overall performance of GLaMM on the target-containing group of BIV-Priv-Seg compared to existing grounding datasets. BIV-Priv-Seg is evaluated in three prompt types, and VizWiz-FewShot is evaluated in the privacy-agnostic scenario only as the dataset contains non-private categories only. The results of GranD, refCOCO, refCOCO+, and refCOCOg are adopted from previous work [39].
Fine-Grained Analysis.
We conducted analyses similar to object size and presence of text as we did for few-shot localization algorithms and additional analysis on object saliency. Because we observed from qualitative results that VLMs tend to segment the salient object in the images, we also evaluated based on whether the target object is the salient object. To do this, we identified whether a target object is salient by applying a salient object detection model [26] that has shown to yield comparable performance to humans for images taken by people with visual impairments (i.e., for the VizWiz dataset) [40] and then applying a $0.5$ threshold to its intersection over union (IoU) with the target object. This means that when over half of the masks are overlapped, the object is considered salient in our analysis. The results of object size and presence of text align with that of few-shot localization, and we found in the results of object saliency that the model performs better on salient objects.
@@
Table 6: Fine-grained analysis of the performance of GLaMM on BIV-Priv-Seg in prompt I scenarios.
Results are presented in Table 6. For object size, we see a consistent trend of performance growth from smaller to larger size; for presence of text, the performance on objects with texts is consistently higher than those on objects without texts. These results correspond to previous findings and reflect the existing gaps between modern algorithms and applications. Concerning object saliency, we observe a strong correlation between saliency and segmentation accuracy. The localization performance is consistently and significantly better on salient objects than on non-salient objects. This relation implies that the modern VLMs rely on an object’s saliency for grounding, as opposed to a complete reliance on the features of the target objects, while also suggesting a potential bias in the existing training data.
@@ 5 Conclusions
While few-shot and zero-shot localization algorithms have made significant strides in recent years, their application to privacy preservation has been limited due to the scarcity of publicly available, annotated images containing private visual information. Yet, such algorithms bring significant potential for building privacy-aware assistive tools because they empower each user or developer to relatively easily embrace distinct definitions of what is private.
To fill this gap, we introduced a new few-shot localization dataset challenge that contributes to the development of privacy-preserving technologies for people who are blind or with low vision: BIV-Priv-Seg. Our extensive analysis of modern algorithms on BIV-Priv-Seg reveals they struggle most with private objects that are small, lack text, and are non-salient. Despite the current gaps between models’ capabilities and real-world privacy preservation needs, the performance of the benchmarked models provides promising prospects for promoting a safer user environment. We offer our work as a valuable foundation for developing more flexible and personalizable privacy-aware solutions that can cater to individual needs.
We envision two important directions for future work. First, is supporting personalized identification of private objects on one’s own devices, where users train models to recognize their private objects in any new images or videos on-device. A critical direction to make this idea a reality is developing such models to be compact so that they can run on resource-constrained devices. A second promising research direction is infusing privacy-aware capabilities into modern VLMs since BLV users already rely on such conversational agents (e.g., Be My AI [13], Microsoft’s Seeing AI [2]). This will enable them to seamlessly realize privacy-preserving capabilities without considerably changing their workflows. Potential interactions include identifying private content in images, obfuscating the information, and providing detailed privacy status updates through a conversational interface. Critical research directions to bring this idea to reality include reconciling VLMs’ inconsistencies between text and mask predictions while also infusing them with the ability to recognize when requested objects are absent from an image.
@@ Acknowledgments.
This project was supported by a National Science Foundation SaTC award (#2148080, #2126314, and #2125925).

### v3

@@ 4.2 VLM Localization
We next benchmarked top-performing, off-the-shelf vision language models that support visual grounding in a zero-shot setting in the following three ways:
Prompt I. Privacy-agnostic (category specific): Each category’s name is inserted into the prompt: “Segment it if there is a {category name} in the given image, otherwise say there is no {category name}."
Prompt II. Privacy-aware (category agnostic): We use the following generic prompt “Segment it if there is a private object in the given image, otherwise say there is no private object.", where we always use the generic phrase “private object". This prompt is intended to capture the model’s understanding regarding what is private.
Prompt III. Privacy-explained: We provide context regarding what is private by providing the list of categories in our dataset in the prompt as follows: “Private objects include {all category names}. Segment it if there is a private object in the given image, otherwise say there is no private object."
We intentionally designed these prompts to support when no target object is present, meaning that the model does not return a segmentation result. We achieved this by extending the prompts from previous work [44] from “Segment something” to “Segment it if something, otherwise say there is no something”.
We tested four models, including three conversational VLMs and one text-prompted segmentation model. However, two of the conversational VLMs failed so terribly that we excluded them from detailed analysis: CogVLM [54] and GPT-4o coupled with Dall-E2 [43]. Specifically, CogVLM achieved a mAP score of 0 with prompts I, II, and III and GPT-4o consistently produced blank binary masks for our prompts and further tested variations (e.g., “Generate a binary mask that segments the {category name}."). We also tested the conversational VLM, GLaMM [44], and the text-prompted segmentation approach, GroundingDINO [33] coupled with Segment Anything (SAM) [29]. Following the implementation in the GroundingDINO [33], the prompts are simplified to single terms for GroundingDINO coupled with Segment Anything (SAM). Specifically, we adopted the designed prompts I and II and prompted the model as follows: (1) {category name}" and (2) “private object".
@@
Privacy classification: For conversational algorithms (e.g., GLaMM), we evaluate the performance on two levels: (1) Semantic level prediction, which is based on the language used in the conversation, where positive responses such as “sure, here is the mask” indicate the presence of a target, and negative responses such as “no, there is no…” indicate the absence of a target; and (2) Mask level prediction, which is determined by whether any pixel in the generated mask has a positive prediction. The results are presented in accuracy, recall, and true negative rate. Accuracy shows the overall performance, the percentage of correct predictions among the target-containing images, and the percentage of correct predictions among the target-less images, respectively. For mask generation-only algorithms (e.g., GroundingDINO coupled with SAM), we only evaluate the results on Mask level prediction.
Segmentation: For the images that are classified as private on both semantic and mask levels, we then perform localization. The mIoU and cIoU (class-weighted IoU) metrics are used to evaluate the semantic segmentation results across all categories in each scenario. As most of the images in this dataset only contain one target object, we convert those segmentation results into instance segmentation format and evaluate the models in AP50 and Recall metrics. This two-step evaluation highlights the model’s ability to identify whether there are private objects in the image and its localization ability.
Privacy Classification Performance.
The results of both the semantic level and mask level privacy classifications are shown in Table 4.
We observe three key findings. First, on the semantic level, the GLaMM consistently generates positive responses across all scenarios, producing masks in every image. However, even then the model confirms the presence of a private object at the conversational level, some of the generated masks do not predict any pixels. This inconsistency suggests the model lacks a robust text-mask alignment mechanism. Second, GLaMM’s negative mask level predictions of are still quite inaccurate given the low true negative (TN) rates in prompt II and III. Finally, GroundingDINO+SAM, unlike GLaMM, does not generate any negative responses. This underscores that models’ struggle to identify when no target objects are in the images. We suspect that these issues stem from biases in the models’ training data, where there is a lack of target-less samples.
@@
Prompt II 81.32 100.00 0.00 69.26 83.97 5.21
Prompt III 81.32 100.00 0.00 63.71 74.28 17.71
GroundingDINO + SAM
Prompt I - - - 81.32 100.00 0.00
Prompt II - - - 81.32 100.00 0.00
Table 4: Privacy classification performance of GLaMM and GroundingDINO coupled with SAM on BIV-Priv-Seg presented in accuracy (Acc), recall (R), and true negative rate (TN) as percentages. Results of GLaMM are evaluated on both semantic and mask levels, and those of GroundingDINO coupled with SAM are evaluated in mask level as there is no textual output.
Evaluated Dataset mIoU cIoU AP50 Recall
GLaMM
BIV-Priv-Seg (ours, prompt I) 68.0 66.9 44.7 55.9
BIV-Priv-Seg (ours, prompt II) 11.4 11.4 3.1 7.8
BIV-Priv-Seg (ours, prompt III) 20.6 20.6 9.4 14.8
VizWiz-FS [52] (prompt I) 15.6 13.4 2.2 5.6
@@
GroundingDINO + SAM
BIV-Priv-Seg (ours, prompt I) 73.2 70.2 60.2 69.6
BIV-Priv-Seg (ours, prompt II) 39.7 33.1 18.8 31.4
VizWiz-FS (prompt I) 39.0 40.2 17.4 35.1
Table 5: Overall performance of the segmentation results of GLaMM and GroundingDINO coupled with SAM on BIV-Priv-Seg. BIV-Priv-Seg is evaluated with all prompts, and VizWiz-FewShot is evaluated with prompt I only as the dataset contains only non-private categories. GLaMM results for GranD, refCOCO, refCOCO+, and refCOCOg are adopted from [44].
Segmentation Performance.
Segmentation results for the target-containing group are shown in Table 5, with qualitative results in Figure 5. We first compared the models’ performance on our dataset to those on VizWiz-FewShot [52], GranD [44], refCOCO, refCOCO+, and refCOCOg [26]. The results show that the performance of our dataset is comparable to that of existing datasets. Specifically, the performance of GLaMM on refCOCO, refCOCO+, and refCOCOg is better than BIV-Priv-Seg (prompt I), while its performance on GranD and VizWiz-FewShot is worse; and GroundingDINO coupled with SAM performs better on BIV-Priv-Seg than VizWiz-FewShot. This indicates that state-of-the-art VLMs have the potential to effectively ground private objects in images, achieving performance comparable to other datasets.
We then compared the performance of different prompts on BIV-Priv-Seg to explore the model’s ability to comprehend privacy. Prompt I achieved the highest performance for both models, indicating that the model has the potential to locate private objects accurately when provided specific object categories. Conversely, prompt II had the worst performance. That may be in part because the models’ notion of what is private may differ from what has been indicated by BLV individuals [51], including because the concept can be highly subjective. Interestingly, prompt III with GLaMM attempted to mitigate this semantic ambiguity by providing an explanation of privacy, yet it still performed poorly. Despite the initial definition of privacy in prompt III, the model failed to fully comprehend and execute the instructions. This suggest the model is highly sensitive to the sentence structure of prompts, which may reflect biases in its training data.
Overall, the performance of GroundingDINO coupled with SAM is better than GLaMM. We hypothesize that the conversational inputs for GLaMM add a challenging complexity for the grounding task, despite it potential benefit of offering end users more flexibility (e.g., prompt III).
We next conducted fine-grained analyses with respect to object size and presence of text, as done inSection 4.1. Also, because we observed from qualitative results that VLMs tend to segment the salient object, we evaluated based on whether the target object is the detected salient object [28] by checking if there was at least a $0.5$ threshold IoU between the segmentations. Results are shown in Table 6.
Results of object size and presence of text show performance increases from smaller to larger objects and yield better performance for objects with text. This reinforces findings from prior work [52] for images taken by BLV individuals, underscoring that certain similarities between private and non-private objects can serve as potential opportunities in overcoming the natural scarcity of private data. We also observe consistently better performance for salient objects than non-salient objects, implying that modern foundation models leverage object saliency for grounding.
@@
GroundingDINO + SAM (prompt I)
mIoU 91.1 67.8 57.3 79.9 91.1 82.5 59.4
cIoU 92.5 65.7 57.3 77.8 92.9 82.1 53.2
AP50 83.3 46.6 35.6 64.5 87.3 69.6 36.8
Recall 84.0 60.2 50.6 67.4 86.5 73.8 53.9
Table 6: Fine-grained analysis of GLaMM and GroundingDINO coupled with SAM on BIV-Priv-Seg in prompt I scenarios.
@@ 5 Conclusions
We introduce a new few-shot localization dataset challenge that contributes to developing privacy-preserving technologies for BLV users: BIV-Priv-Seg. While there remain gaps between models’ capabilities and privacy preservation needs, our public-sharing of the dataset offers a valuable foundation for future advancement of privacy-aware solutions. Our results from model benchmarking underscore that a promising direction for future research is infusing privacy-aware capabilities into modern VLMs, particularly because BLV users already rely on such conversational agents (e.g., Be My AI [14], Microsoft’s Seeing AI [2]). More generally, we expect success on our dataset will facilitate progress for a broader number of scenarios facing similar challenges, such as for robotics and lifelogging.
@@ Acknowledgments.
This project was supported by National Science Foundation SaTC awards (#2148080, #2126314, and #2125925).